Circular No. 8873

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 S1 (ZHAO) Haibin Zhao, Purple Mountain Observatory, reports the discovery of

a comet (discovery observation tabulated below) on CCD images obtained with the 1.04-m f/1.8 Schmidt telescope at XuYi Station; the object shows a moderately condensed coma with diameter 15′′.6. Following posting on the ‘NEOCP’ webpage, G. Sostero (Remanzacco, Italy) reports that unfiltered CCD images obtained by L. Donato and V. Gonano on Sept. 19.9 UT show (0.45-m reflector; forty co-added 60-s exposures) a coma nearly 25′′ in diameter with a fan-shaped tail nearly 40′′ long in p.a. 240^{o}. L. Buzzi and F. Luppi (Varese, Italy, 0.60-m reflector) write that their forty-five stacked 15-s images from Sept. 20.0 show a 10′′ central condensation elongated in p.a. ∼ 270^{o} surrounded by a faint 30′′ coma. E. Guido, Castellammare di Stabia, Italy, writes that 12 stacked 60-s unfiltered exposures taken remotely by Sostero and himself on Sept. 20.4 with a 0.25-m f/3.4 reflector at Mayhill, NM, U.S.A., show a diffuse coma nearly 15′′ in diameter with total mag 17.8 and a sharp central condensation with a tail nearly 25′′ long toward p.a. 260^{o}. Ten 180-s images taken by R. Ligustri (Talmassons, Udine, Italy, 0.35-m reflector) on Sept. 20.9 show a coma of diameter ∼ 20′′–30′′.

2007 UT  α_{2000}  δ_{2000}  Mag.
Sept.17.78693  0^{h}47^{m}37.^{s}51  −2^{o}00′44′′.2  18.2

The available astrometry (including prediscovery LONEOS observations from Sept. 8 and 15 identified by T. Spahr), the following elliptical orbital elements, and an ephemeris appear on MPEC 2007-S36.

T = 2007 Dec. 10.559 TT  ω = 246.^{o}072
e = 0.34443  Ω = 141.930 }2000.0
q = 2.50287 AU  i = 6.024
a = 3.81784 AU n^{o} = 0.132123 P = 7.46 years

SATELLITES OF SATURN The IAU Working Group for Planetary System Nomenclature has ap-

proved new designations and names of satellites of Saturn (cf. IAUC 8826):

Saturn XLIX Anthe  = S/2007 S 4  IAUC 8857
Saturn L  Jarnsaxa = S/2006 S 6  IAUC 8727
Saturn LI  Greip  = S/2006 S 4  IAUC 8727
Saturn LII  Tarqeq = S/2007 S 1  IAUC 8836
2007 September 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green